Can someone help me with this question: A rectangle animal pen will be built using 200 meters of fencing. If one side of the rectangle is 60meters, find the area of the pen.
Found 2 solutions by stanbon, Alan3354:
Answer by stanbon(75887)

A rectangle animal pen will be built using 200 meters of fencing. If one side of the rectangle is 60meters, find the area of the pen.
-----
Draw the rectangle.
Label one side as 60 meters.
---
Then the side opposite is also 60 meters.
Label both of the other sides as "x":
Equation:
Perimeter = (each of the 4 sides)
200 = 2*60+2x
---
100 = 60 + x
x = 40 meters.
---
Area = 40*60 = 2400 sq. ft.
=============================
Cheers,
Stan H.
=============================

Answer by Alan3354(69443)

Can someone help me with this question: A rectangle [sic] animal pen will be built using 200 meters of fencing. If one side of the rectangle is 60meters, find the area of the pen.
-----------
Hint:
Perimeter = 2(L + W) = 200
L + W = 100
----
Area = L*W